Nightlog 20171228 UT (E/D \xBD, AZ \xBD)

Cast of Characters:

OSA: David Gonzalez Huerta (mountain)

Isp: Alessandro Cavallaro (mountain)

Observer: Olga Kuhn (remote room), joined by Andrew Rakich and John Hill (by phone) during

the LBC TMS observing.

Plan:

Start with LBCR, TMS testing, if seeing is <1.2\x94 switch to MODS ~9pm for collection of transform data. If seeing > 1.2\x94, switch to MODS for D targets (FF0240), or quasars from poor condition and not-so-poor conditions samples.

00:17 UT Opened.

00:55 TMS laser was running all day and we\x92re watching it track M1 position/size as

temperatures change. Tamb/Tglass = 4.6/4.4 and ambient from envgui = 3.9.

01:00 Slew to field at ~71 deg elevation for collimation: 01:30:00 30:00:00 (close to M33)

01:00 dohybrid, /redonly \x85 done

01:13 Andrew setting his reference, I run lbc_set_tms_offset.py

01:16 Taking an in-focus image at r\x92 011732

01:19 Run tms_loop.py. Not running any LBC OB.. We see the instrument offset (0.025 mic for

r) but then see a jump. Letting it converge...Stopping

01:22 Another in-focus image. 012226. Holes in the middles - out of focus. (loop pushed it

away\x85)

01:25 Run dofpia.

01:33 lbc_set_tms_offset.py (referenced at -0.8mm)

01:36 in-focus image 013717

01:38 starting tms_loop.py (see -0.8mm plus 0.25). Jumped again. Andrew calling John Hill.

01:39 taking another image to see whether we have donuts. Yes, see 013953.

01:47 slew to new field (2017 WE2) with tms_loop running. Lots of \x91nan\x92s; Andrew has a

beam-break.

01:51 -- \x93found a bug in our algorithm\x94. Need to record offsets before apply z11 (loop script

applies z11).

01:56 John going to turn off z11 in loop script --- editing script.

01:56 dofpia, /redonly.

02:03 lbc_set_tms_offset

02:06 in-focus 5-sec image 020722

02:08 start tms_loop. No jump now, but there is a large tilt --- loop not driving down to 0:

0,125,~0,rx is largish, (this is the pointing change).

02:12 take another image with loop still running (02:14 - tms software glitch)

02:15 dofpia, /redonly elev~55 deg.

02:19 set_offsets

02:20 superfoc 7 @ r

02:27 infocus image 022745

02:28 r_centextra 023003

02:31 tms_loop \x85 once converged we\x92ll go back to superfoc. But although it was going to 0, we

then got into a Y limit, -2.95mm. Polynomial correction? Group rotation Y= -5.613.

02:33 Pressed collimate but still at Y limit. Cntrl\x92Cd loop (did not stop\x85. Depends when you

control c) and removed group rotation corrections. That brought it out of a limit. Finally

control-c\x92d.

02:38 Remove group rot corrections. Starting tms_loop. Hit limit again. (tms_loop -l will print

more info). John and Andrew debugging. We did not clear group rotation before running dofpia.

02:44 tracking same field\x85 remove group rotation

02:44 run dofpia, /redonly

02:50 Andrew set ref, we run set_offset, then tms_loop to insure it is not hitting limits. Not hitting

limits, but some discussion about pointing corrections\x85 if fpia is done at low elevation, we\x92d hit

limits.

Starting series with superfocs\x85

1:

02:58 in-focus image 025858

03:00 superfoc started (moonlit clouds\x85but moving)

03:06 RB_rcentextra

03:08 tms_loop\x85 until group rotation numbers stable, then cntl-c

2:

03:10 superfoc

03:15 in-foc image

03:18 rcentextra

3:

03:20 Tms_loop\x85

03:23 superfoc

03:30 tms_loop, infocus image 033054

03:34 tms_loop, rcentextra

Slewing to a new field. We want a field that will transit in about 1 hr. This field will reach 80 deg

elevation. We\x92ll defocus at 9:20pm (20 min before focus).

03:44 tms_loop on new field (expect largeish change in Y, RX\x85 seen\x85)

03:50 infocus image 035041

03:54 tms_loop

03:54 superfoc

04:00 tms_loop

04:01 centextra

04:02 tms_loop...

04:04 in-foc image 040519

Tms_loop

04:09 superfoc

04:19 tms_loop, rcentextra

04:22 Now we\x92d like to maintain a certain defocus, of -0.35mm. John will edit

/tmp/tms_right_20171228_lbcobs. file to make offset 0.45 and not 0.8. This will mimic the

defocus of -0.35mm.

04:25 run tms_loop

04:29 starting series of 15-sec exposures on this field with defocus -0.35mm. Will continue this

through transit which will be in ~9 min to get a series of defocussed images through transit. We

can difference these\x85 we don\x92t see growing coma, but a bit of astigmatism that is changing in

orientation. Tms corrects coma and focus, not astigmatism or spherical.

05:00 Going to switch to MODS for collection of transform data. LBC TMS got good data.

05:25 cp PCSInstrument_correctMODSzps.conf PCSInstrument.conf

05:29 David bouncing PCS

05:30 First will do acq on StoneD1 - after pointing/collim

05:50 acqBinoMODS mods.49.StoneD1f.acq

mods1Align:

Dpa, dx, dy = 0.032, -0.027, -0.24

mods2Align:

Dpa, dx, dy = 0.0453, 1.562, 0.842

OK\x85

06:05 iif_register, ldg and iif_register_rdg

06:06 David slewing to StoneD star 1 PA=270.. Problem with preset (maybe I need to send it?)

06:14 meanwhile, ~modseng/modsScripts/Support/transformInit.pro

06:21 try sending preset from idl following twiki instructions

iif_preset,Stone_D,mode=active,gs=1,ts=1,pa=270,/right (/left)

06:47 collecting transform data. David could preset to Stone_D but not at rotator angle 270,

which is the PA for which I\x92d made the lists.

Found other lists at PA0 and doing these.

Transform_collect \x85 Stone_D_star17_pa0.list.

Taking a mods image on blue side\x85 OK (dual imaging mode being used).

Mods2 b 2 image with probe on-axis. Transform_collect for left side is not working.

MODS2 is taking blue images, not red ones\x85 OK, but science lam = 5650 so it is a little closer

to r. We should do r. Maybe we had a problem with MODS2r the last time we did this so

switched to MODS2b.

I can run transform_collect but we hit cp limit immediately. Try right only.

07:06 David authorizing right side only.

07:08: starting transform_collect on Stone_D_star7_pa270.list. MODS2 only.

At this point we are not going to get much, but we have a short time to try.

Will probably try to do standard for AZ program since it is clear now - more valuable than

collecting a few stars only. (Well, first acq for AZ should start at 1am\x85 will see how transform

progresses through list --- maybe enough data to determine sign of rotation??)

Well --- still there are error messages

IRCR: preset, Stone_D, MODS=GUIDE, GS=7, TS=7, PA=270.000 (iif_preset)

⇒ guiding stopped

⇒ GS setting didn\x92t succeed (this is for on-axis)

⇒ isPosAccessible: probe cannot move to requested position (SFP) (x,y): 0,150mm.

⇒ problem moving the AGW probe onto the target position

⇒ OK

Continues along anyway\x85

5 has probe on axis and is supposedly BG field

6 will be gs=0, ts=7 (first measurement)

I\x92ll make up scripts to do the MODS transform data collection --- it will be simpler, I think. As is,

we are taking an unnecessary background image and I always forget how to run the iif_ scripts,

plus we have some unexpected issues, like why I cannot send iif_preset command, but can

send presets via the transform_collect\x85 must be user error\x85 sigh!!!

Let this run through to do the 12 stars.

7 gs=1

8 gs=2


active

9 gs=3

10 gs=4

11 gs=6


active

12 gs=8

13 gs=9

14 gs=10


active

15 gs=15

17 gs=21 (last one)


active (why?)

Transform_mods2.log in /home/lbto/data/20171228/rdg

(need to fix image numbering)

07:49 Restore PCSInstrument.conf file, bounce PCS and authorize both

-- %USERSIG{OlgaKuhn - 2017-12-29}%

Comments


 
Topic revision: r1 - 29 Dec 2017, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback